Tech Staff Summary 7/10/00

Current Priorities in addition to routine tasks

Prison at Tiptonville needs another 2 days work (to be completed 7/11-12).
Covington and Troy and done. Troy needs a phone line (is Bob arranging?).
Reelfoot is more or less complete.
Steele is done.
One antenna upgrade is left to do.
PI's are searching for the hard rock site to the east.

PARM will need a new borehole and will be reinstalled the week of Jul 17. BRGM will be fixed week of Jun 17.

Greg and Steve are organizing mockup of the comlete WAN.
Greg has sent agreements to the 5 tower owners and will check status.
Greg has bid's in hand and is working with Purchasing.
Mitch will make sure we have 2 laptops for the installation to check signal strengths.
We can begin installation of equipment at towers that aren't subcontracted anytime (august).
Greg is delevoping a scheme for getting the data from the unnamed tower into the computer room without going through the CERI router in the long building.
Greg will need to visit each site and coordinate with owner/engineers.
Greg has bids for antenna installation.

PPTN is up but still needs some tweaking.

Interagency agreement with TN Forest Service is on the way.
Copper Ridge has a 20' soil cap; should we make it ISISII?
MTTN node will be installed when there are 3 or 4 stations up (September?). Hussey needs a 3 week notice.
John is here this week to get the old suburban to replace his truck.
Mitch is working on insurance claim.
Mitch is still waiting to hear from Shedlock before determining target location for the CMG-3.
Knoxville/TVA node is still slated for September. Mitch is working with Williams and Benz to get this organized.
Steve is looking into possibilities for getting ATHN continuously.
LRNC has extremely poor telemetry and marginal internet bandwidth. Possible fix for the telemetry is reactivation of High Peak as a repeater. Greg will get with John about this.

Board Shop
On hold, preferred location is in the dark room of the long building.
When we find out where it will be we will either install a door providing access to the darkroom from the expansion space or perform the necessary OSHA upgrades to the existing location.

Jim and Chris were cruising until other projects took priority. Need two more weeks to complete LNXT and NMAD which we can hopefully squeeze in June/July/August.
Mitch still needs to complete the code to handle the gain channel. Subroutines are done and offline PII style is available. On the fly will need less robust than offline to avoid potentially large and time intensive buffering.
Problems exist with some upgraded stations. Mitch will provide some appropriate earthquake data to determine the problems and aid in the solution.

On schedule for 7/24 installation. We will digitize the IRIGH and stamp with pc system time. Running xntp will help keep this accurate. Consider replacing ISDN with DSL.
Greg is developing a scheme to allow pumping I-40 bridge data through here if necessary.

Given resolution of current plans, the current neotectonics lab will become the geophysical observatory space.
An outside cage will be constructed for battery charging for general CERI needs.
Appropriate cable and equipment racks will be constructed and installed in the neotectonics lab.
PC's and workstations will go where Tucker is now.
Dirty equipment will go into Chris's garage.
Needs to be operational before September arrival of PASSCAL equipment.

Fish Fry was a great success. We should consider doing this more often.

Second tier priorities

Cannabalize PI to make as much as PII operational.
Check with Arch or Jer-Ming for account numbers to buy supplies.
Not sure where the space exists to work on this.

EBZ/PWLA telemetry. Another repeater?

LNXT node Secure propane tank.
Secure racks.
replace coax jumpers
replace rg-58 unshielded jumpers

MKTA node

NMAD node

Need to record the SOH channel from the gps clocks to know when sync is lost (they will continue to produce IRIGE based on internal clock).

Backburner priorities.

Station book. Mitch and Steve will install earthworm v5 in June. This will include the full blown schema capable of creating seed.

Status channels. Currently supplied on last 8 channels.

LNXT node permit/install L7 and L8

MKTA node permit/install remaining 3 ISIS

NMAD node